# Multimodal Medical Image Binding via Shared Text Embeddings

### v1

@@ Abstract.
Medical image analysis increasingly relies on the integration of multiple imaging modalities to capture complementary anatomical and functional information, enabling more accurate diagnosis and treatment planning. Achieving aligned feature representations across these diverse modalities is therefore important for effective multimodal analysis. While contrastive language-image pre-training (CLIP) and its variant have enabled image-text alignments, they require explicitly paired data between arbitrary two modalities, which is difficult to acquire in medical contexts. To address the gap, we present Multimodal Medical Image Binding with Text (M3Bind), a novel pre-training framework that enables seamless alignment of multiple medical imaging modalities through a shared text representation space without requiring explicit paired data between any two medical image modalities. Specifically, based on the insight that different images can naturally bind with text, M3Bind first fine-tunes pre-trained CLIP-like image-text models to align their modality-specific text embedding space while preserving their original image-text alignments. Subsequently, we distill these modality-specific text encoders into a unified model, creating a shared text embedding space. Experiments on X-ray, CT, retina, ECG, and pathological images on multiple downstream tasks demonstrate that M3Bind achieves state-of-the-art performance in zero-shot, few-shot classification and cross-modal retrieval tasks compared to its CLIP-like counterparts. These results validate M3Bind’s effectiveness in achieving cross-image-modal alignment for medical analysis.
Multimodal Medical Image Analysis, Contrastive Learning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, lrelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligence
@@
Figure 1. We illustrate the M3Bind framework, designed to embed multiple medical imaging modalities (X-ray, CT, retina, ECG, and Pathology) within a cohesive text-anchored space. In this figure, the transition from Phase One to Phase Two represents the five modality-specific CLIP-like models we used. These models start by aligning each modality with the text modality in isolation and then use text as the anchor modality, achieving joint alignment of all modalities through alignment with different text encoders. The progression from Phase Two to Phase Three indicates the further distillation of knowledge from the five aligned modality-specific text encoders into a unified text encoder, to achieve a consistent representation of all texts paired with different modalities.
1. Introduction
@@
Medical diagnosis is increasingly based on the integration of multiple imaging modalities that provide complementary anatomical and pathological information (Hussain et al., 2022). For example, computed tomography (CT) enables high resolution 3D tissue visualization (Rubin, 2014), chest radiographs offer a broad structural context (Speets et al., 2006), retina imaging captures microvascular patterns (Liew et al., 2008), electrocardiogram (ECG) records the electrical activity of the heart (Mirvis and Goldberger, 2001), and histopathology provides tissue visualization at the cell level (Majno and Joris, 2004). While each modality contributes unique diagnostic information, their full potential emerges when analyzed together in a multimodal fashion (Davnall et al., 2012). For example, the evaluation of respiratory diseases benefits from integrating the detailed visualization of the lung parenchyma of CT with the broader anatomical context provided by X-rays (Chicklore et al., 2013). Similarly, cardiovascular diagnostics leverage the synergy between retinal vasculature patterns and histopathological tissue examination to evaluate both structural and functional aspects of cardiac health (O’connor et al., 2017). Conditions such as pericarditis require the combination of ECG data showing changes in the ST segment and PR segment with chest radiographs that reveal the cardiac silhouette or pericardial calcification (Adler et al., 2015). These clinical scenarios highlight the importance of aligned feature representations that can effectively synthesize diagnostic information across diverse medical imaging modalities (Glynne-Jones et al., 2014).
@@
Contrastive language image pre-training (CLIP) (Radford et al., 2021) can align embedding spaces between visual and textual information through contrastive learning (Li et al., 2022a, b; Mu et al., 2022; Yao et al., 2021). In medical image analysis, specialized CLIP variants have been proposed to integrate specific medical images with the corresponding clinical text descriptions (Zhao et al., 2023). For radiology applications, GLoRIA (Huang et al., 2021) and MGCA (Wang et al., 2022) improve diagnostic interpretability by aligning chest X-rays with radiology reports, while CT-CLIP (Hamamci et al., 2024) extends this framework to volumetric CT data analysis. In ophthalmology and pathology domains, FLAIR (Silva-Rodriguez et al., 2025) and QuiltNet (Ikezogwo et al., 2023) respectively leverage semantic supervision for zero-shot retinal diagnosis and fine-grained histological interpretation. However, these approaches typically focus on alignment between a single medical imaging modality and text, limiting their applicability to multiple medical imaging modalities (Wang et al., 2022; Huang et al., 2021). For $n$ different modalities, conventional CLIP-based approaches would require training (${n\choose 2}=\frac{n(n-1)}{2}$ separate alignment models, which is not only computationally intensive but also limited by the difficulty of acquiring the corresponding paired data between different imaging types.
@@
Recent work in multimodal learning demonstrates that diverse modalities can be aligned in a shared embedding space without requiring explicit pairwise training. Specifically, ImageBind (Girdhar et al., 2023) uses natural images as a bridge to connect six different data modalities through their inherent connections to visual information. However, direct application of ImageBind to medical imaging is not applicable due to the lack of an anchor visual modality. Alternatively, “binding through images” concept inspires our approach to multimodal medical imaging, where diverse modalities frequently co-occur with clinical text descriptions that can serve as semantic bridges. For example, radiological images are routinely accompanied by detailed reports, pathology slides with diagnostic annotations, and retina scans with clinical evaluations. This motivates the use of clinical text as a semantic anchor to align diverse medical imaging modalities in a unified representation space (Figure 1).
@@
To bridge the gap, we propose multimodal medical image binding with text (M3Bind), a multimodal pre-training framework for medical image analysis. M3Bind innovates by using text as an anchor modality to bind to various medical imaging modalities. Unlike previous CLIP-like methods, M3Bind aligns different image modalities through a shared text representation space without requiring explicit paired data between any two image modalities. We further employ distillation (Hinton, 2015; Jiao et al., 2019; Mirzadeh et al., 2020; Sanh, 2019; Touvron et al., 2021; Wu et al., 2022) to consolidate modality-specific text encoders into a unified model that enables seamless cross-modal analysis.
@@
The main contributions are four-fold. Firstly, we propose M3Bind, a novel multimodal pre-training framework that addresses the challenge of cross-modal medical image alignment by using text as a central bridging modality, enabling seamless integration of multiple CLIP-like image-text encoders without requiring explicit paired data between image modalities. Secondly, we propose an adaptive training strategy that effectively counters the inherent modality imbalance between different medical image modalities, ensuring balanced representation learning across diverse modalities. Thirdly, we introduce Shared Embedding Space Knowledge Distillation to consolidate representations from multiple modality-specific text encoders into a single unified text model, enhancing computational efficiency while preserving alignment quality across modalities. Finally, we conduct a comprehensive evaluation of M3Bind on 12 diverse datasets that span X-ray, CT, retina, ECG, and pathological images, demonstrating state-of-the-art performance in zero shot classification, few shot classification, and cross-modal retrieval tasks compared to existing CLIP-based approaches.
2. Related Works
2.1. CLIP Variants
@@
CLIP has inspired numerous adaptations that improve multimodal representation learning in diverse domains. Early attempts focused on representation quality, where SLIP (Mu et al., 2022) incorporated self-supervised learning objectives alongside contrastive training to improve feature robustness, while DeCLIP (Smeu et al., 2024) leveraged multi-dimensional supervision signals to increase data efficiency. Several later approaches have explored masking strategies to optimize the alignment process. For example, MaskCLIP (Dong et al., 2023) employed masked self-distillation with an exponential moving average (EMA) encoder to strengthen alignment quality and noise resilience, while FLIP (Li et al., 2023) and Alpha-CLIP (Sun et al., 2024) introduced input masking that balances computational efficiency with model depth and performance. Most relevant to our work, ImageBind (Girdhar et al., 2023) advanced multimodal integration by unifying six distinct modalities, including images, text, audio, depth, thermal, and IMU data, into an aligned representation space using images as a bridge modality. Although ImageBind demonstrated impressive performance in tasks in the natural domain, its approach cannot be directly applied to medical imaging contexts where no single image modality can serve as a universal anchor.
2.2. Medical CLIP-like Models
@@
Medical imaging has seen adaptation of CLIP to address domain-specific challenges, which leverage contrastive learning to align medical images with clinical text. For volumetric data analysis, CT-CLIP (Hamamci et al., 2024) aligns 3D computed tomography scans with radiology reports. In ultrasound imaging, EchoCLIP (Christensen et al., 2024) connects echocardiogram video frames with clinical descriptions to improve cardiac function assessment. Ophthalmology applications benefit from FLAIR (Silva-Rodriguez et al., 2025), which enriches the analysis of retinal images through the supervision of semantic text, substantially improving zero shot diagnostic capabilities. For chest radiography, MGCA (Wang et al., 2022) proposes multi-granularity cross-modal contrastive learning to capture hierarchical disease manifestations. In digital pathology, QuiltNet (Ikezogwo et al., 2023) integrates histopathological images with expert knowledge for the interpretation of fine-grained tissue. Despite these advances, each model operates within a single imaging modality, limiting their utility in comprehensive multimodal analysis scenarios. MedBind (Gao et al., 2024) represents an initial step toward multimodal integration in medicine by introducing a specialized loss function to handle paired X-ray and ECG data. However, MedBind still relies on explicitly curated image-to-image pairings, which are difficult to obtain at scale across diverse medical imaging modalities. This dependency on paired data presents a constraint for broader multimodal medical image analysis. Our proposed M3Bind addresses this limitation by introducing a text-centric binding approach that eliminates the need for direct image-to-image pairings.
@@
Figure 2. Overview of the M3Bind. We start with five modality-specific CLIP-like models (X-ray, CT, retina, ECG, and pathology). The M3Bind integrates these models by maintaining image-text alignment within each modality while aligning text representations across different modalities. LoRA is applied to efficiently fine-tune these pre-trained models while preserving their original capabilities. The Shared Embedding Space Knowledge Distillation (SESKD) consolidates knowledge from all modality-specific text encoders (teachers) into a unified BioClinicalBERT-based student model. This student model is trained using both MSE loss for direct embedding alignment and contrastive loss to preserve semantic relationships with all five imaging modalities.
2.3. Knowledge Distillation
@@
In single-modality contexts, Knowledge Distillation (KD) achieved notable success by compressing large networks while preserving performance through the transfer of soft probability distributions rather than hard labels(Hinton, 2015). Regarding multimodal fields, recent attempts include DistillVLM (Wang et al., 2023), which introduced cross-modal attention alignment to preserve vision-language interactions between teacher and student models. Additionally, TinyCLIP (Wu et al., 2023) used affinity mimicking and weight inheritance to optimize vision language embeddings, although its effectiveness remains constrained by architectural similarity requirements. CLIP-KD (Yang et al., 2024a) overcomes this limitation by supporting knowledge transfer between heterogeneous architectures. Further advances have refined the granularity of knowledge transfer in multimodal models. DIME-FM (Sun et al., 2023) proposed instance-level distillation to capture fine-grained relationships between individual samples, while CLIP-CID (Yang et al., 2024b) extended this approach with cluster-level distillation that preserves semantic relationships between instance groups. In our M3Bind, we leverage KD to address the challenge of unifying multiple modality-specific text encoders into a single model while preserving alignment with diverse medical imaging modalities.
3. Methods
3.1. Method Overview
@@
3.2. M3Bind for Pre-training
M3Bind introduces a novel pertaining method that leverages text as a universal anchor to align five distinct medical imaging modalities (X-ray, CT, retina, ECG, and pathology) without requiring paired image data between image modalities. Specifically, M3Bind starts with CLIP-like medical models, each pre-trained on a specific medical imaging modality (X-ray, CT, retina, ECG, and pathology) with corresponding textual descriptions. Building upon five CLIP variants namely MGCA (Wang et al., 2022), CT-CLIP (Hamamci et al., 2024), FLAIR (Silva-Rodriguez et al., 2025), ECG-CLIP (proposed by MedBind(Gao et al., 2024) and trained by ourselves), and QuiltNet (Ikezogwo et al., 2023), M3Bind begins with these modality-specific pre-trained image-text models, each trained on their respective medical imaging modality with corresponding textual descriptions. The first stage of M3Bind establishes text as a shared semantic space by aligning the text embedding spaces of these five CLIP-like models while preserving their respective original image-text alignments through contrastive learning. Specifically, for each modality $m$, we maintain the alignment of image and text using the following contrastive loss:
(1) $\mathcal{L}_{CLIP}^{(m)}=-\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}\left(\mathbf{I}_{i}^{(m)},\mathbf{T}_{i}^{(m)}\right)/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}\left(\mathbf{I}_{i}^{(m)},\mathbf{T}_{j}^{(m)}\right)/\tau\right)},$
@@ 3.2. M3Bind for Pre-training
(2) $\mathcal{L}_{MSE}^{(m_{1},m_{2})}=\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{T}_{m_{1},i}-\mathbf{T}_{m_{2},i}\right|^{2}$
@@
3.3. Adaptive Modality Balancing

### v2

@@ Abstract
Medical image analysis increasingly relies on the integration of multiple imaging modalities to capture complementary anatomical and functional information, enabling more accurate diagnosis and treatment planning. Achieving aligned feature representations across these diverse modalities is therefore important for effective multimodal analysis. While contrastive language-image pre-training (CLIP) and its variant have enabled image-text alignments, they require explicitly paired data between arbitrary two modalities, which is difficult to acquire in medical contexts. To address the gap, we present Multimodal Medical Image Binding with Text (M3Bind), a novel pre-training framework that enables seamless alignment of multiple medical imaging modalities through a shared text representation space without requiring explicit paired data between any two medical image modalities. Specifically, based on the insight that different images can naturally bind with text, M3Bind first fine-tunes pre-trained CLIP-like image-text models, which are derived from different medical modalities, to align their modality-specific text embedding space while preserving their original image-text alignments. Subsequently, we distill these modality-specific text encoders into a unified model, creating a shared text embedding space. Notably, M3Bind is a flexible framework in which the selection of CLIP-like models is not fixed and can be adapted according to the requirements of the task. Experiments on X-ray, CT, retina, ECG, and pathological images on multiple downstream tasks demonstrate that M3Bind achieves competitive or even superior performance in zero-shot, few-shot classification and cross-modal retrieval tasks compared to its CLIP-like counterparts. These results validate M3Bind’s effectiveness in achieving cross-image-modal alignment for medical analysis.
1 Introduction
@@
Figure 1: We illustrate the M3Bind framework, designed to embed multiple medical imaging modalities (X-ray, CT, retina, ECG, and Pathology) within a cohesive text-anchored space. In this figure, the transition from Phase One to Phase Two represents the five modality-specific CLIP-like models we used. These models start by aligning each modality with the text modality in isolation and then use text as the anchor modality, achieving joint alignment of all modalities through alignment with different text encoders. The progression from Phase Two to Phase Three indicates the further distillation of knowledge from the five aligned modality-specific text encoders into a unified text encoder, to achieve a consistent representation of all texts paired with different modalities.
@@
Medical diagnosis is increasingly based on the integration of multiple imaging modalities that provide complementary anatomical and pathological information [18]. For example, computed tomography (CT) enables high resolution 3D tissue visualization [35], chest radiographs offer a broad structural context [40], retina imaging captures microvascular patterns [27], electrocardiogram (ECG) records the electrical activity of the heart [30], and histopathology provides tissue visualization at the cell level [29]. While each modality contributes unique diagnostic information, their full potential emerges when analyzed together in a multimodal fashion [5]. For example, the evaluation of respiratory diseases benefits from integrating the detailed visualization of the lung parenchyma of CT with the broader anatomical context provided by X-rays [4]. Similarly, cardiovascular diagnostics leverage the synergy between retinal vasculature patterns and histopathological tissue examination to evaluate both structural and functional aspects of cardiac health [33]. Conditions such as pericarditis require the combination of ECG data showing changes in the ST segment and PR segment with chest radiographs that reveal the cardiac silhouette or pericardial calcification [1]. These clinical scenarios highlight the importance of aligned feature representations that can effectively synthesize diagnostic information across diverse medical imaging modalities [11].
@@
Contrastive language image pre-training (CLIP) [34] can align embedding spaces between visual and textual information through contrastive learning [24, 25, 32, 52]. In medical image analysis, specialized CLIP variants have been proposed to integrate specific medical images with the corresponding clinical text descriptions [56]. For radiology applications, GLoRIA [16] and MGCA [46] improve diagnostic interpretability by aligning chest X-rays with radiology reports, while CT-CLIP [13] extends this framework to volumetric CT data analysis. In ophthalmology and pathology domains, FLAIR [38] and QuiltNet [19] respectively leverage semantic supervision for zero-shot retinal diagnosis and fine-grained histological interpretation. However, these approaches typically focus on alignment between a single medical imaging modality and text, limiting their applicability to multiple medical imaging modalities [46, 16]. For $n$ different modalities, conventional CLIP-based approaches would require training ${n\choose 2}=\frac{n(n-1)}{2}$ separate alignment models, which is not only computationally intensive but also limited by the difficulty of acquiring the corresponding paired data between different imaging types.
@@
Recent work in multimodal learning demonstrates that diverse modalities can be aligned in a shared embedding space without requiring explicit pairwise training. Specifically, ImageBind [10] uses natural images as a bridge to connect six different data modalities through their inherent connections to visual information. However, direct application of ImageBind to medical imaging is not applicable due to the lack of an anchor visual modality. Alternatively, “binding through images” concept inspires our approach to multimodal medical imaging, where diverse modalities frequently co-occur with clinical text descriptions that can serve as semantic bridges. For example, radiological images are routinely accompanied by detailed reports, pathology slides with diagnostic annotations, and retina scans with clinical evaluations. This motivates the use of clinical text as a semantic anchor to align diverse medical imaging modalities in a unified representation space (Figure 1).
@@
To bridge the gap, we propose multimodal medical image binding with text (M3Bind), a multimodal pre-training framework for medical image analysis. M3Bind innovates by its fine-tuning of CLIP-like models across multiple modalities, using text as an anchor modality to bind to various medical imaging modalities. Unlike previous CLIP-like methods, M3Bind creates a shared text representation space by aligning the text modalities of all CLIP-like models, thereby aligning different image modalities without requiring explicit paired data between any two image modalities. We further employ distillation [14, 21, 31, 36, 43, 48] to consolidate modality-specific text encoders into a unified model that enables seamless cross-modal analysis.
@@
The main contributions are four-fold. Firstly, we propose M3Bind, a novel multimodal pre-training framework that addresses the challenge of cross-modal medical image alignment by using text as a central bridging modality, enabling seamless and flexible integration of multiple CLIP-like image-text encoders without requiring explicit paired data between image modalities. Secondly, we propose an adaptive training strategy that effectively counters the inherent modality imbalance between different medical image modalities, ensuring balanced representation learning across diverse modalities. Thirdly, we introduce Shared Embedding Space Knowledge Distillation to consolidate representations from multiple modality-specific text encoders into a single unified text model, enhancing computational efficiency while preserving alignment quality across modalities. Finally, we conduct a comprehensive evaluation of M3Bind on 12 diverse datasets that span X-ray, CT, retina, ECG, and pathological images, demonstrating state-of-the-art performance in zero shot classification, few shot classification, and cross-modal retrieval tasks compared to existing CLIP-based approaches.
2 Related Work
2.1 CLIP Variants
@@
CLIP has inspired numerous adaptations that improve multimodal representation learning in diverse domains. Early attempts focused on representation quality, where SLIP [32] incorporated self-supervised learning objectives alongside contrastive training to improve feature robustness, while DeCLIP [39] leveraged multi-dimensional supervision signals to increase data efficiency. Several later approaches have explored masking strategies to optimize the alignment process. For example, MaskCLIP [7] employed masked self-distillation with an exponential moving average (EMA) encoder to strengthen alignment quality and noise resilience, while FLIP [26] and Alpha-CLIP [42] introduced input masking that balances computational efficiency with model depth and performance. Most relevant to our work, ImageBind [10] advanced multimodal integration by unifying six distinct modalities, including images, text, audio, depth, thermal, and IMU data, into an aligned representation space using images as a bridge modality. Although ImageBind demonstrated impressive performance in tasks in the natural domain, its approach cannot be directly applied to medical imaging contexts where no single image modality can serve as a universal anchor.
2.2 Medical CLIP-like Models
@@
Medical imaging has seen adaptation of CLIP to address domain-specific challenges, which leverage contrastive learning to align medical images with clinical text. For volumetric data analysis, CT-CLIP [13] aligns 3D computed tomography scans with radiology reports. Ophthalmology applications benefit from FLAIR [38], which enriches the analysis of retinal images through the supervision of semantic text, substantially improving zero shot diagnostic capabilities. For chest radiography, MGCA [46] proposes multi-granularity cross-modal contrastive learning to capture hierarchical disease manifestations. In digital pathology, QuiltNet [19] integrates histopathological images with expert knowledge for the interpretation of fine-grained tissue. Despite these advances, each model operates within a single imaging modality, limiting their utility in comprehensive multimodal analysis scenarios. BiomedCLIP[53] and UniMed-CLIP[22] aim to collect more diverse and extensive data to train models. However, these methods employ a single shared visual encoder for images across different modalities. Given the significant heterogeneity among multimodal medical images, this approach inevitably limits the representational capacity of the model. MedBind [9] represents an initial step toward multimodal integration in medicine by introducing a specialized loss function to handle paired X-ray and ECG data. However, MedBind still relies on explicitly curated image-to-image pairings, which are difficult to obtain at scale across diverse medical imaging modalities. This dependency on paired data presents a constraint for broader multimodal medical image analysis. Our proposed M3Bind addresses this limitation by introducing a text-centric binding approach that eliminates the need for direct image-to-image pairings.
@@
Figure 2: Overview of the M3Bind. We start with five modality-specific CLIP-like models (X-ray, CT, retina, ECG, and pathology). The M3Bind integrates these models by maintaining image-text alignment within each modality while aligning text representations across different modalities. LoRA is applied to efficiently fine-tune these pre-trained models while preserving their original capabilities. The Shared Embedding Space Knowledge Distillation (SESKD) consolidates knowledge from all modality-specific text encoders (teachers) into a unified BioClinicalBERT-based student model. This student model is trained using both MSE loss for direct embedding alignment and contrastive loss to preserve semantic relationships with all five imaging modalities.
2.3 Knowledge Distillation
@@
In single-modality contexts, Knowledge Distillation (KD) achieved notable success by compressing large networks while preserving performance through the transfer of soft probability distributions rather than hard labels[14]. Regarding multimodal fields, recent attempts include DistillVLM [47], which introduced cross-modal attention alignment to preserve vision-language interactions between teacher and student models. Additionally, TinyCLIP [49] used affinity mimicking and weight inheritance to optimize vision language embeddings, although its effectiveness remains constrained by architectural similarity requirements. CLIP-KD [50] overcomes this limitation by supporting knowledge transfer between heterogeneous architectures. Further advances have refined the granularity of knowledge transfer in multimodal models. DIME-FM [41] proposed instance-level distillation to capture fine-grained relationships between individual samples, while CLIP-CID [51] extended this approach with cluster-level distillation that preserves semantic relationships between instance groups. In our M3Bind, we leverage KD to address the challenge of unifying multiple modality-specific text encoders into a single model while preserving alignment with diverse medical imaging modalities.
3 Methods
3.1 Method Overview
@@
3.2 M3Bind for Pre-training
M3Bind introduces a novel pertaining method that leverages text as a universal anchor to align five distinct medical imaging modalities (X-ray, CT, retina, ECG, and pathology) without requiring paired image data between image modalities. Specifically, M3Bind starts with CLIP-like medical models, each pre-trained on a specific medical imaging modality (X-ray, CT, retina, ECG, and pathology) with corresponding textual descriptions. Building upon five CLIP variants namely MGCA [46], CT-CLIP [13], FLAIR [38], ECG-CLIP (proposed by MedBind[9] and trained by ourselves), and QuiltNet [19], M3Bind begins with these modality-specific pre-trained image-text models, each trained on their respective medical imaging modality with corresponding textual descriptions. The selection of these models is flexible and can be freely chosen according to specific needs, not limited to the models we have actually selected. The first stage of M3Bind establishes text as a shared semantic space by aligning the text embedding spaces of these five CLIP-like models while preserving their respective original image-text alignments through contrastive learning. Specifically, for each modality $m$, we maintain the alignment of image and text using the following contrastive loss:
$\mathcal{L}_{CLIP}^{(m)}=-\sum_{i=1}^{N}\log\frac{\exp\left(\text{sim}\left(\mathbf{I}_{i}^{(m)},\mathbf{T}_{i}^{(m)}\right)/\tau\right)}{\sum_{j=1}^{N}\exp\left(\text{sim}\left(\mathbf{I}_{i}^{(m)},\mathbf{T}_{j}^{(m)}\right)/\tau\right)},$ (1)
@@ 3.2 M3Bind for Pre-training
$\mathcal{L}_{MSE}^{(m_{1},m_{2})}=\frac{1}{N}\sum_{i=1}^{N}\left|\mathbf{T}_{m_{1},i}-\mathbf{T}_{m_{2},i}\right|^{2}$ (2)
@@
3.3 Adaptive Modality Balancing